Determine Ek. of water with volume V = 3m3, if it moves with speed v = 72km / h.


1
Expert's answer
2021-01-14T12:51:38-0500

By definition, the kinetic energy is:


"E_k =\dfrac{mv^2}{2}"

where "m" is the mass of the water, and "v = 72km\/h = 20m\/s" is its speed. The mass of "V = 3m^3" is:


"m = 1000\dfrac{kg}{m^3}V = 3000kg"

Thus, obtain:


"E_k =\dfrac{3000\cdot 20^2}{2} = 0.6\times 10^{6}J = 0.6MJ"

Answer.  0.6 MJ.